# DoctorAgent-RL: A Multi-Agent Collaborative Reinforcement Learning System for Multi-Turn Clinical Dialogue

### v1

@@ Abstract
Large language models (LLMs) have demonstrated excellent capabilities in the field of biomedical question answering, but their application in real-world clinical consultations still faces core challenges. Existing systems rely on a one-way information transmission mode where patients must fully describe their symptoms in a single round, leading to nonspecific diagnostic recommendations when complaints are vague. Traditional multi-turn dialogue methods based on supervised learning are constrained by static data-driven paradigms, lacking generalizability and struggling to intelligently extract key clinical information. To address these limitations, we propose DoctorAgent-RL, a reinforcement learning (RL)-based multi-agent collaborative framework that models medical consultations as a dynamic decision-making process under uncertainty. The doctor agent continuously optimizes its questioning strategy within the RL framework through multi-turn interactions with the patient agent, dynamically adjusting its information-gathering path based on comprehensive rewards from the Consultation Evaluator. This RL fine-tuning mechanism enables LLMs to autonomously develop interaction strategies aligned with clinical reasoning logic, rather than superficially imitating patterns in existing dialogue data. Notably, we constructed MTMedDialog, the first English multi-turn medical consultation dataset capable of simulating patient interactions. Experiments demonstrate that DoctorAgent-RL outperforms existing models in both multi-turn reasoning capability and final diagnostic performance, demonstrating practical value in assisting clinical consultations.
1 Introduction
Large language models (LLMs) such as ChatGPT achiam2023gpt , LLaMA touvron2023llama , and ChatGLM glm2024chatglm have demonstrated remarkable capabilities in various natural language processing tasks, including open-domain question answering, dialogue generation, and code synthesis chen2025benchmarking . With their strong generalization abilities, these models are increasingly applied to healthcare domains, where they exhibit potential for providing preliminary medical advice and assisting clinical decision-making kopka2025accuracy .
The core flaw of current medical LLMs lies in the inherent contradiction between their passive single-turn interaction paradigm and the active information-gathering requirements of clinical diagnosis. Although existing systems demonstrate a high level of knowledge accuracy in structured medical question-answering tasks, their "answer-only" characteristic forces patients to fully describe their symptoms in a single round of conversation wu2025towards . However, in reality, complex medical conditions often require multiple targeted inquiries to gradually clarify the situation laban2025llms . This misalignment in interaction patterns leads to two systemic limitations: On the one hand, users tend to provide vague or fragmented initial descriptions, and the diagnostic suggestions generated by the models based on incomplete information often lack specificity and may even pose potential risks auerbach2024diagnostic . On the other hand, existing methods that attempt to introduce multi-turn dialogue capabilities through supervised learning, limited by the pattern imitation of static dialogue datasets, fail to establish a dynamic decision-making mechanism to weigh the value of information against dialogue efficiency liu2025interactive .
@@ 1 Introduction
To address these challenges, we propose DoctorAgent-RL—a multi-agent collaborative reinforcement learning framework that reformulates clinical reasoning as a Markov Decision Process (MDP). Within this framework: (1) A high-fidelity patient agent based on LLMs, which generates pathologically consistent responses while mimicking the diversity of real-world communication; (2) A clinician agent initialized by cloning medical actions from real consultation records and refined through reinforcement learning (RL) to master effective questioning strategies; and (3) A consultation evaluator that provides multi-dimensional rewards based on diagnostic accuracy, patient information responsiveness, and the standardization of questions. Notably, we constructed MTMedDialog, the first English multi-turn medical consultation dataset capable of simulating realistic patient interactions. This dataset enables the training and evaluation of multi-agent systems in dynamic clinical reasoning scenarios. DoctorAgent-RL redefines the diagnostic process through reward-based strategic policy optimization: The doctor agent continuously optimizes its questioning strategy within a RL framework, obtaining immediate feedback through multi-turn interactions with the patient agent, and dynamically adjusting the information-gathering path based on the comprehensive rewards provided by the consultation evaluator. This end-to-end RL fine-tuning mechanism enables large language models to autonomously evolve interaction strategies that align with clinical reasoning logic, rather than simply imitating surface patterns in existing dialogue data. Experiments demonstrate that DoctorAgent-RL outperforms existing models in both multi-turn reasoning capability and final diagnostic performance, as shown in Figure  1, demonstrating practical value in assisting clinical consultations. Through such multi-agent collaborative training, DoctorAgent-RL allows the model to essentially learn how to ask context-relevant questions and when to confidently terminate the consultation, thereby achieving proactive and strategic patient interaction. Our contributions are three-fold:
@@
  • We propose DoctorAgent-RL, a multi-agent collaborative RL framework where the doctor agent autonomously develops clinically-aligned questioning strategies through interactions with the patient agent, guided by the consultation evaluator’s comprehensive reward mechanism. DoctorAgent-RL dynamically adjusts information-gathering pathways, overcoming the limitations of conventional static data-driven paradigms.
@@
  • DoctorAgent-RL achieves state-of-the-art performance in multi-turn medical consultations, demonstrating exceptional capabilities in generating high-quality questions and delivering more accurate clinical diagnoses.
Figure 1: Comparison of the average accuracy scores for diagnosis and recommendation, as well as the average interaction rounds, between DoctorAgent-RL and other models.
2 Related Work
2.1 Medical QA Systems
Despite achieving high accuracy in standardized benchmark tests, current medical LLMs are still restricted by their passive single-turn QA pattern tu2024towards . Systems like MedAlpaca han2023medalpaca and MedDialog zeng2020meddialog concentrate on optimizing single-response generation via supervised fine-tuning, with BioMistral labrak2024biomistral , a medical adaptation of Mistral mistral7b , demonstrating particularly strong few-shot performance on clinical benchmarks, which implicitly presumes that users can offer comprehensive symptom descriptions. Nevertheless, this assumption contradicts real-world scenarios where patients typically initial with vague complaints. Although models such as HuaTuo wang2023huatuo and UltraMedical zhang2024ultramedical enhance answer quality by incorporating medical knowledge graphs, they still function reactively, compelling patients to serve as self-diagnosticians by interpreting general advice like “consult a doctor”. This limitation arises from framing medical QA as a language generation task rather than a sequential decision process. Current approaches prioritize maximizing token-level prediction accuracy but fail to model the clinician’s key challenge of strategically eliciting critical information through adaptive questioning nori2024medprompt . Thus, even state-of-the-art systems struggle with ambiguous cases where diagnosis relies on iterative hypothesis refinement, a capability beyond single-turn interaction design.
2.2 Multi-turn Dialogue Systems
@@
Although systems like Bianque chen2023bianque and DialoGPT zhang2019dialogpt attempt to introduce multi-turn capabilities through supervised fine-tuning on synthetic or annotated dialogue datasets, they remain constrained by static training paradigms. The standardized multiple-choice evaluation framework liao2023automatic aims to mitigate LLMs’ hallucination risks through structured testing, yet its scripted key-value dialogue protocol fails to capture authentic doctor-patient interactions. Recent approaches like the APP system zhu2025ask propose entropy-minimized diagnostic optimization, achieving higher symptom recall than conventional methods through medical guideline grounding. MDDial macherla2023mddial contributes a specialized dataset of 1,200 expert-annotated diagnostic dialogues with turn-level reliability scores, though its template-based generation limits linguistic diversity and complexity, failing to fully reflect the richness of real clinical conversations. These methods still rely on predefined dialogue paths rather than learning optimal questioning strategies from clinical outcomes. While systems incorporating turn reliability scoring and iterative differential diagnosis surpass single-turn response capabilities, they fundamentally suffer from static training data limitations and lack mechanisms to optimize dialogue strategies based on interaction results li2024mediq . Consequently, their ability to ask the right questions at the right time remains inadequate.
2.3 Multi-Agent Systems in Medicine
@@
Recent advances in medical multi-agent systems have demonstrated significant potential in clinical reasoning simulation and diagnostic accuracy improvement, yet fundamental limitations persist in achieving human-centric adaptability. AMIE tu2025towards optimize diagnostic dialogues through self-play simulation environments, outperforming primary care doctors in structured evaluations, but their reliance on static training data restricts adaptability to novel clinical scenarios. The MAC framework chen2025enhancing enhances rare disease diagnosis by simulating multidisciplinary team discussions, yet its template-driven dialogue patterns lack linguistic diversity and contextual flexibility. Agent Hospital li2024agent incorporate dynamic medical record libraries and LLM-generated disease progression models to evolve agent behaviors, though the generated experiences carry risks of factual inconsistencies. While the AI Hospital framework fan2024ai improves diagnostic accuracy through multi-role collaboration protocols, it fails to adequately account for the diversity of treatment options and the effectiveness of alternative strategies in real clinical settings. Current systems neither achieve dynamic integration of real-time medical evidence nor adjust questioning depth based on patient comprehension levels, resulting in suboptimal information delivery efficiency.
2.4 Reinforcement Learning in Medicine
@@
The application of RL in the medical field is transitioning from static decision-making to dynamic interaction paradigms, demonstrating unique value in optimizing strategies for multi-turn consultation systems. MedVLM-R1 pan2025medvlm and Med-R1 lai2025med employ the Group Relative Policy Optimization (GRPO) shao2024deepseekmath framework, which integrates RL reward mechanisms with radiological imaging feature analysis to incentivize vision-language models in generating interpretable reasoning pathways. This approach significantly enhances diagnostic accuracy and reduces hallucinatory reasoning, yet exhibits limited adaptability to emerging modalities. HuatuoGPT-o1 chen2024huatuogpt enhances clinical reasoning through verifiable question generation and medical validation feedback mechanisms, but its reliance on multiple-choice question data conversion limits adaptability to unstructured symptom descriptions. MedRIA zou2024ai employs actor-critic frameworks to optimize inquiry efficiency in emergency scenarios, yet requires manual feature engineering for complex diagnoses. PPME sun2025improving leverages clinical experience replay to prioritize high-value diagnostic pathways, achieving domain-specific accuracy gains at the cost of poor generalization. These studies collectively indicate that RL requires further refinement in dynamic interaction objective modeling and clinically oriented evaluation frameworks to address the complexity and uncertainty inherent in real-world medical practice.
3 Method
3.1 Dataset and Evaluation Metrics
3.1.1 Dataset creation process
@@
We present MTMedDialog, the first English multi-turn medical consultation dataset capable of simulating patient interactions. The dataset comprises 8,086 training samples and 2,082 test samples derived from three Chinese benchmark datasets: IMCS21 10.1093/bioinformatics/btac817 , CHIP-MDCFNPC zhang2021cblue , and MedDG liu2022meddg . The test set covers 8 major disease categories, with detailed distribution statistics provided Appendix B.
@@
As the source datasets were collected from real doctor-patient conversations, we implemented a two-stage denoising strategy: (1) filtering shallow dialogues with less than three turns through exact turn-count matching, and (2) removing noisy segments containing consecutive meaningless responses using DeepSeek-V3 liu2024deepseek . We strictly preserved the original data partitioning protocol during this cleaning process to ensure evaluation reliability. The retained dialogues were then translated into English using DeepSeek-V3.
@@
3.1.2 Evaluation metrics
The doctor agent evaluation focuses on two core dimensions: (1) Diagnosis and Recommendation Accuracy: This dimension evaluates eight disease categories using a 6-level quantitative matching scale (0=completely incorrect, 5=exact match). Semantic consistency between the agent’s diagnosis/recommendations and gold-standard labels is assessed via Qwen2.5-32B-Instruct xu2025qwen2 , with specific prompts shown in Appendix C. The final score is converted to a 100-point scale by multiplying the raw score by 20; (2) Interaction Turns: The average number of dialogue turns required for the doctor agent and the patient agent to reach a final diagnosis.
The patient agent assessment employs a three-dimensional DeepSeek-V3-based scoring system: (1) Information Control: Quantifies unsolicited information disclosure through a baseline score of 1.0 (perfect compliance), with 0.2-point deductions per unauthorized disclosure instance; (2) Response Completeness: Evaluates critical information omission against doctor queries using an initial score of 1.0, penalized by 0.2 points per essential detail omission; (3) Factual Conflict: Detects contradictions with medical records through a violation counter starting at 0.0, accumulating 0.2 points per identified inconsistency. Patient agent’s rating prompts shown in Appendix C.
3.2 Task Formulation
@@ 3.2.1 Doctor Agent
The reward signal $R\in\mathbb{R}$ for each trajectory is provided by an independent consultation evaluator (detailed in Section 3.2.3), considering diagnostic accuracy, information acquisition efficiency, and protocol compliance as metrics.
To enhance the stability of policy optimization and eliminate the requirement for an additional value function approximation, we propose Group Relative Policy Optimization (GRPO) shao2024deepseekmath as our policy gradient algorithm. Unlike Proximal Policy Optimization (PPO) schulman2017proximal , GRPO employs the average reward of multiple sampled outputs as a baseline instead of relying on a learned value function. Specifically, for each patient’s consultation $x$, GRPO samples a set of trajectories ${y_{1},y_{2},\dots,y_{G}}$ through the interaction between the doctor agent $\pi_{D}$ and the patient agent $\pi_{\text{p}}$. The doctor agent, as the policy model, is then optimized by maximizing the following objective function:
$\mathcal{J}_{GRPO}(\theta)=\,$ $\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\text{D}_{\text{old}}}(\cdot|x;\pi_{\text{p}})}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\sum_{t=1}^{|y_{i}|}I(y_{i,t})}\sum_{t=1:I(y_{i,t})=1}^{|y_{i}|}\min\Bigg{(}\frac{\pi_{\text{D}}(y_{i,t}|x,y_{i,<t})}{\pi_{\text{D}_{\text{old}}}(y_{i,t}|x,y_{i,<t})}\hat{A}_{i,t},$
$\hskip 120.0pt\text{clip}\Bigg{(}\frac{\pi_{\text{D}}(y_{i,t}|x,y_{i,<t})}{\pi_{\text{D}_{\text{old}}}(y_{i,t}|x,y_{i,<t})},1-\epsilon,1+\epsilon\Bigg{)}\hat{A}_{i,t}\Bigg{)}-\beta\mathbb{D}_{KL}\left[\pi_{\text{D}}||\pi_{\text{D}_{ref}}\right]\Bigg{]},$ (1)
@@
3.2.2 Patient Agent
The patient agent is implemented using Qwen2.5-7B-Instruct xu2025qwen2 and is incorporated into a two-phase dialogue simulation framework via carefully crafted prompt engineering.
@@ 3.2.2 Patient Agent
In the second phase, the patient agent utilizes dynamic symptom release strategies in response to the real-time queries of the doctor agent. It maintains strict pathological consistency while mimicking the natural variability in patients’ symptom description granularity and the order of their complaints. The detailed prompt designs for both phases are presented in Appendix E. By retaining complete hidden case data, the patient agent ensures that its natural language responses adhere to clinical standards and are generated solely based on the dialogue history.
3.2.3 Consultation Evaluator
@@
Information Acquisition Efficiency Reward. A proficient clinician knows how to ask the right questions efficiently. To foster this skill in our doctor agent, the Consultation Evaluator incorporates a dynamic reward mechanism that promotes valuable questioning while discouraging repetitive or unhelpful queries. This reward is directly tied to the patient agent’s feedback after each dialogue turn and accumulates throughout the interaction:
@@
$R_{\text{compliance}}^{t}=\begin{cases}-2,&\text{if the question format violates predefined norms}\\ -5,&\text{if no diagnosis is provided within the allowed turns}\\ 0,&\text{otherwise (i.e., protocol is followed)}\end{cases}$ (5)
@@
This sophisticated, multi-dimensional Consultation Evaluator not only guides the model toward superior diagnostic accuracy but also actively encourages the development of efficient information-gathering strategies and professional, compliant clinical behavior, ultimately aiming to achieve diagnostic capabilities that closely resemble those of expert human clinicians.
3.3 Training Procedure
As illustrated in Figure 2, our training framework for the doctor agent is built upon Qwen2.5-7B-Instruct, following the DeepSeek-R1 training paradigm guo2025deepseek . The approach employs a two-stage training pipeline, integrating SFT and RL to cultivate clinical reasoning capabilities. Detailed experimental settings are illustrated in Appendix H.
@@ 3.3 Training Procedure
Following SFT, we refine the agent’s decision-making under interaction constraints using the policy optimization algorithm detailed in Section 3.2. To enhance robustness and mimic real-world clinical scenarios, we introduce a Dynamic Turn Budget Training Strategy. Each training episode is assigned a random dialogue turn budget (2–10 turns). The model is explicitly reminded of the remaining turns after each interaction step, encouraging efficient information gathering. This two-stage approach ensures the agent first internalizes clinical reasoning patterns via supervised learning, then refines its strategy through interactive reward optimization. Ablation studies comparing alternative training strategies (direct SFT, direct RL, fixed-turn training) are presented in the experimental section.
4 Experiments
4.1 Patient Agent Behavior Analysis on MTMedDialog
Table  1 presents the evaluation results of the effectiveness of patient agent in simulating real-world medical interactions. In the dimension of information control, Qwen2-5.7B demonstrates optimal compliance performance, precisely constraining output content and effectively avoiding the leakage of irrelevant information. In terms of response completeness, while DeepSeek-V3 performs best, Qwen2-5.7B still maintains near-optimal levels of key information density required for doctor-patient dialogue. Notably, all tested models achieve zero error rates in the dimension of factual conflicts, fully validating the reliability of LLMs in medical knowledge. Considering both model performance and cost-effectiveness in training, we ultimately selected Qwen2-5.7B as the implementation solution for the patient agent. For a more detailed description of the dialogue flow, please refer to Appendix G

### v2

@@ Abstract
Large language models (LLMs) have demonstrated excellent capabilities in the field of biomedical question answering, but their application in real-world clinical consultations still faces core challenges. Single-round consultation systems require patients to describe all symptoms upfront, leading to vague diagnosis with unclear complaints. Traditional multi-turn dialogue models, constrained by static supervised learning, lack flexibility and fail to intelligently extract key clinical information. To address these limitations, we propose DoctorAgent-RL, a reinforcement learning (RL)-based multi-agent collaborative framework that models medical consultations as a dynamic decision-making process under uncertainty. The doctor agent continuously optimizes its questioning strategy within the RL framework through multi-turn interactions with the patient agent, dynamically adjusting its information-gathering path based on comprehensive rewards from the Consultation Evaluator. This RL fine-tuning mechanism enables LLMs to autonomously develop interaction strategies aligned with clinical reasoning logic, rather than superficially imitating patterns in existing dialogue data. Notably, we constructed MTMedDialog, the first English multi-turn medical consultation dataset capable of simulating patient interactions. Experiments demonstrate that DoctorAgent-RL outperforms existing models in both multi-turn reasoning capability and final diagnostic performance. This approach shows immense practical value by reducing misdiagnosis risks in time-pressured settings, freeing clinicians for complex cases, and pioneering a strategy to optimize medical resource allocation and alleviate workforce shortages.
Introduction
Large language models (LLMs) like ChatGPT (Achiam et al. 2023), LLaMA (Touvron et al. 2023), and ChatGLM (GLM et al. 2024) show significant potential for assisting in clinical decision-making (Kopka, von Kalckreuth, and Feufel 2025). This capability is urgently needed in the context of a global doctor shortage and high-pressure healthcare environments (Agyeman-Manu et al. 2023). In these settings, time pressure on clinicians often leads to rapid diagnoses based on incomplete information, increasing misdiagnosis risks (ALQahtani et al. 2016), while routine consultations consume resources that could be allocated to more critical cases (Arndt et al. 2024). Therefore, developing an intelligent system to improve primary care efficiency holds significant practical importance for optimizing resource allocation and alleviating healthcare strain (Al Kuwaiti et al. 2023).
Existing systems suffer from two fundamental limitations. First, the ”answer-only” characteristic of single-turn systems (Wu et al. 2025) requires patients to describe their condition comprehensively at once, which conflicts with the clinical reality that complex conditions often necessitate multi-turn inquiries for clarification (Laban et al. 2025). This misalignment can lead to nonspecific or even risky diagnostic suggestions based on incomplete information (Auerbach et al. 2024). Second, current multi-turn dialogue methods that rely on fine-tuning are constrained by imitating static data patterns, and thus lack the dynamic decision-making mechanism required to reason like a clinician (Liu et al. 2025).
Figure 1: Overview of this work. DoctorAgent-RL achieves state-of-the-art on the MTMedDialog task.
@@ Introduction
To address these challenges, we propose DoctorAgent-RL—a multi-agent collaborative reinforcement learning framework that reformulates clinical reasoning as a Markov Decision Process (MDP), as shown in Figure 1. Within this framework: (1) A high-fidelity patient agent based on LLMs, which generates pathologically consistent responses while mimicking the diversity of real-world communication; (2) A doctor agent initialized by cloning medical actions from real consultation records and refined through reinforcement learning (RL) to master effective questioning strategies; and (3) A consultation evaluator that provides multi-dimensional rewards based on diagnostic accuracy, patient information responsiveness, and the standardization of questions. Notably, we constructed MTMedDialog, the first English multi-turn medical consultation dataset capable of simulating realistic patient interactions. This dataset enables the training and evaluation of multi-agent systems in dynamic clinical reasoning scenarios. DoctorAgent-RL redefines the diagnostic process through reward-based strategic policy optimization: The doctor agent continuously optimizes its questioning strategy within a RL framework, obtaining immediate feedback through multi-turn interactions with the patient agent, and dynamically adjusting the information-gathering path based on the comprehensive rewards provided by the consultation evaluator. This end-to-end RL fine-tuning mechanism enables LLMs to autonomously evolve interaction strategies that align with clinical reasoning logic, rather than simply imitating surface patterns in existing dialogue data. Experiments demonstrate that DoctorAgent-RL outperforms existing models in both multi-turn reasoning capability and final diagnostic performance, demonstrating practical value in assisting clinical consultations. This work demonstrates a clear pathway for developing intelligent systems that can assist in primary consultations, thereby reducing misdiagnosis risk, freeing clinicians for complex cases, and helping to optimize healthcare resource allocation.
Our contributions are three-fold:
@@
  • We propose DoctorAgent-RL, a multi-agent RL framework where a doctor agent learns questioning strategies via reward-driven interactions with a patient agent, guided by a consultation evaluator.
@@
  • DoctorAgent-RL achieves state-of-the-art performance in multi-turn medical consultations, excelling in both strategic question generation and final diagnostic accuracy.
Related Work
Medical QA Systems
Current medical LLMs are primarily limited by a passive, single-turn QA pattern (Tu et al. 2024). Systems like MedAlpaca (Han et al. 2023), MedDialog (Zeng et al. 2020), and the Mistral-adapted BioMistral (Labrak et al. 2024) focus on single-response generation, which incorrectly assumes users can provide comprehensive initial symptom descriptions. Even models enhanced with knowledge graphs, such as HuaTuo (Wang et al. 2023) and UltraMedical (Zhang et al. 2024), remain reactive, often providing generic advice that shifts the diagnostic burden back to the patient. This core limitation stems from framing medical QA as a language generation task rather than a sequential decision process, thus failing to model the strategic, adaptive questioning needed for ambiguous cases (Nori et al. 2024).
Multi-turn Dialogue Systems
@@
Although systems like Bianque (Chen et al. 2023) and DialoGPT (Zhang et al. 2019) attempt to introduce multi-turn capabilities through supervised fine-tuning on synthetic or annotated dialogue datasets, they remain constrained by static training paradigms. The standardized multiple-choice evaluation framework (Liao et al. 2023) aims to mitigate LLMs’ hallucination risks through structured testing, yet its scripted key-value dialogue protocol fails to capture authentic doctor-patient interactions. Recent approaches like the APP system (Zhu and Wu 2025) propose entropy-minimized diagnostic optimization, achieving higher symptom recall than conventional methods through medical guideline grounding. These methods still rely on predefined dialogue paths rather than learning optimal questioning strategies from clinical outcomes. While systems incorporating turn reliability scoring and iterative differential diagnosis surpass single-turn response capabilities, they fundamentally suffer from static training data limitations and lack mechanisms to optimize dialogue strategies based on interaction results (Li et al. 2024b). Consequently, their ability to ask the right questions at the right time remains inadequate.
Multi-Agent Systems in Medicine
@@
Recent advances in medical multi-agent systems have demonstrated significant potential in clinical reasoning simulation and diagnostic accuracy improvement, yet fundamental limitations persist in achieving human-centric adaptability. AMIE (Tu et al. 2025) optimize diagnostic dialogues through self-play simulation environments, outperforming primary care doctors in structured evaluations, but their reliance on static training data restricts adaptability to novel clinical scenarios. The MAC framework (Chen et al. 2025) enhances rare disease diagnosis by simulating multidisciplinary team discussions, yet its template-driven dialogue patterns lack linguistic diversity and contextual flexibility. Agent Hospital (Li et al. 2024a) incorporate dynamic medical record libraries and LLM-generated disease progression models to evolve agent behaviors, though the generated experiences carry risks of factual inconsistencies. While the AI Hospital framework (Fan et al. 2024) improves diagnostic accuracy through multi-role collaboration protocols, it fails to adequately account for the diversity of treatment options and the effectiveness of alternative strategies in real clinical settings. Current systems neither achieve dynamic integration of real-time medical evidence nor adjust questioning depth based on patient comprehension levels, resulting in suboptimal information delivery efficiency.
Reinforcement Learning in Medicine
@@
The application of RL in the medical field is transitioning from static decision-making to dynamic interaction paradigms, demonstrating unique value in optimizing strategies for multi-turn consultation systems. MedVLM-R1 (Pan et al. 2025) and Med-R1 (Lai et al. 2025) employ the Group Relative Policy Optimization (GRPO) (Shao et al. 2024) framework, which integrates RL reward mechanisms with radiological imaging feature analysis to incentivize vision-language models in generating interpretable reasoning pathways. This approach significantly enhances diagnostic accuracy and reduces hallucinatory reasoning, yet exhibits limited adaptability to emerging modalities. HuatuoGPT-o1 (Chen et al. 2024) enhances clinical reasoning through verifiable question generation and medical validation feedback mechanisms, but its reliance on multiple-choice question data conversion limits adaptability to unstructured symptom descriptions. MedRIA (Zou et al. 2024) employs actor-critic frameworks to optimize inquiry efficiency in emergency scenarios, yet requires manual feature engineering for complex diagnoses. PPME (Sun et al. 2025) leverages clinical experience replay to prioritize high-value diagnostic pathways, achieving domain-specific accuracy gains at the cost of poor generalization. These studies collectively indicate that RL requires further refinement in dynamic interaction objective modeling and clinically oriented evaluation frameworks to address the complexity and uncertainty inherent in real-world medical practice.
Methodology
Dataset and Evaluation Metrics
Dataset creation process
@@
We present MTMedDialog, the first English multi-turn medical consultation dataset capable of simulating patient interactions. The dataset comprises 8,086 training samples and 2,082 test samples derived from three Chinese benchmark datasets: IMCS21 (Chen et al. 2022), CHIP-MDCFNPC (Zhang et al. 2021), and MedDG (Liu et al. 2022). The test set covers 8 major disease categories, with detailed distribution statistics provided in Supplementary Material.
@@
As the source datasets were collected from real doctor-patient conversations, we implemented a two-stage denoising strategy: (1) filtering shallow dialogues with less than three turns through exact turn-count matching, and (2) removing noisy segments containing consecutive meaningless responses using DeepSeek-V3 (Liu et al. 2024). We strictly preserved the original data partitioning protocol during this cleaning process to ensure evaluation reliability. The retained dialogues were then translated into English using DeepSeek-V3. Finally, we performed a manual inspection and refinement of the test set.
@@
Evaluation metrics
The doctor agent evaluation focuses on two core dimensions: (1) Diagnosis and Recommendation Accuracy: This dimension evaluates eight disease categories using a 6-level quantitative matching scale (0=completely incorrect, 5=exact match). Semantic consistency between the agent’s diagnosis/recommendations and gold-standard labels is assessed via Qwen2.5-32B-Instruct (Xu et al. 2025), with specific prompts shown in Supplementary Material. The final score is converted to a 100-point scale by multiplying the raw score by 20; (2) Interaction Turns: The average number of dialogue turns required for the doctor agent and the patient agent to reach a final diagnosis.
The patient agent assessment employs a three-dimensional DeepSeek-V3-based scoring system: (1) Information Control: Quantifies unsolicited information disclosure through a baseline score of 1.0 (perfect compliance), with 0.2-point deductions per unauthorized disclosure instance; (2) Response Completeness: Evaluates critical information omission against doctor queries using an initial score of 1.0, penalized by 0.2 points per essential detail omission; (3) Factual Conflict: Detects contradictions with medical records through a violation counter starting at 0.0, accumulating 0.2 points per identified inconsistency. Patient agent’s rating prompts shown in Supplementary Material.
Task Formulation
@@
The reward signal $R\in\mathbb{R}$ for each trajectory is provided by an independent consultation evaluator (detailed in Section Consultation Evaluator), considering diagnostic accuracy, information acquisition efficiency, and protocol compliance as metrics.
To enhance the stability of policy optimization and eliminate the requirement for an additional value function approximation, we propose Group Relative Policy Optimization (GRPO) (Shao et al. 2024) as our policy gradient algorithm. Unlike Proximal Policy Optimization (PPO) (Schulman et al. 2017), GRPO employs the average reward of multiple sampled outputs as a baseline instead of relying on a learned value function. Specifically, for each patient’s consultation $x$, GRPO samples a set of trajectories ${y_{1},y_{2},\dots,y_{G}}$ through the interaction between the doctor agent $\pi_{D}$ and the patient agent $\pi_{\text{p}}$. The doctor agent, as the policy model, is then optimized by maximizing the following objective function:
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}_{\begin{subarray}{c}x\sim\mathcal{D}\\ \{y_{i}\}_{i=1}^{G}\sim\pi_{\mathrm{D_{old}}}(\cdot|x;\pi_{\mathrm{p}})\end{subarray}}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\sum_{t=1}^{|y_{i}|}I(y_{i,t})}\sum_{\begin{subarray}{c}t=1\\ I(y_{i,t})=1\end{subarray}}^{|y_{i}|}\\ \min\Bigg{(}\frac{\pi_{\mathrm{D}}(y_{i,t}|x,y_{i,<t})}{\pi_{\mathrm{D_{old}}}(y_{i,t}|x,y_{i,<t})}\hat{A}_{i,t},\ \text{clip}\Bigg{(}\frac{\pi_{\mathrm{D}}(y_{i,t}|x,y_{i,<t})}{\pi_{\mathrm{D_{old}}}(y_{i,t}|x,y_{i,<t})},1-\epsilon,1+\epsilon\Bigg{)}\hat{A}_{i,t}\Bigg{)}\\ -\beta\mathbb{D}_{\mathrm{KL}}\left[\pi_{\mathrm{D}}||\pi_{\mathrm{D_{ref}}}\right]\Bigg{]}$ (1)
@@
Patient Agent
The patient agent is implemented using Qwen2.5-7B-Instruct (Xu et al. 2025) and is incorporated into a two-phase dialogue simulation framework via carefully crafted prompt engineering.
@@
In the second phase, the patient agent utilizes dynamic symptom release strategies in response to the real-time queries of the doctor agent. It maintains strict pathological consistency while mimicking the natural variability in patients’ symptom description granularity and the order of their complaints. The detailed prompt designs for both phases are presented in Supplementary Material. By retaining complete hidden case data, the patient agent ensures that its natural language responses adhere to clinical standards and are generated solely based on the dialogue history.
Consultation Evaluator
@@
Information Acquisition Efficiency Reward. A proficient doctor knows how to ask the right questions efficiently. To foster this skill in our doctor agent, the Consultation Evaluator incorporates a dynamic reward mechanism that promotes valuable questioning while discouraging repetitive or unhelpful queries. This reward is directly tied to the patient agent’s feedback after each dialogue turn and accumulates throughout the interaction:
@@
$R_{\text{compliance}}^{t}=\begin{cases}-2,&\text{Question Format Violation}\\ -5,&\text{No Diagnosis within Allowed Turns}\\ 0,&\text{otherwise (i.e., protocol is followed)}\end{cases}$ (5)
@@
This sophisticated, multi-dimensional Consultation Evaluator not only guides the model toward superior diagnostic accuracy but also actively encourages the development of efficient information-gathering strategies and professional, compliant clinical behavior, ultimately aiming to achieve diagnostic capabilities that closely resemble those of expert human doctors.
Training Procedure
As illustrated in Figure 2, our training framework for the doctor agent is built upon Qwen2.5-7B-Instruct, following the DeepSeek-R1 training paradigm (Guo et al. 2025). The approach employs a two-stage training pipeline, integrating SFT and RL to cultivate clinical reasoning capabilities. Detailed experimental settings are illustrated in Supplementary Material.
@@
Following SFT, we refine the agent’s decision-making under interaction constraints using the policy optimization algorithm detailed in Section Task Formulation. To enhance robustness and mimic real-world clinical scenarios, we introduce a Dynamic Turn Budget Training Strategy. Each training episode is assigned a random dialogue turn budget (2–10 turns). The model is explicitly reminded of the remaining turns after each interaction step, encouraging efficient information gathering. This two-stage approach ensures the agent first internalizes clinical reasoning patterns via supervised learning, then refines its strategy through interactive reward optimization. Ablation studies comparing alternative training strategies (direct SFT, direct RL, fixed-turn training) are presented in the experimental section.
Experiment
Patient Agent Behavior Analysis on MTMedDialog
Table  1 presents the evaluation results of the effectiveness of patient agent in simulating real-world medical interactions. In the dimension of information control, Qwen2.5-7B demonstrates optimal compliance performance, precisely constraining output content and effectively avoiding the leakage of irrelevant information. In terms of response completeness, while DeepSeek-V3 performs best, Qwen2.5-7B still maintains near-optimal levels of key information density required for doctor-patient dialogue. Notably, all tested models achieve zero error rates in the dimension of factual conflicts, fully validating the reliability of LLMs in medical knowledge. Considering both model performance and cost-effectiveness in training, we ultimately selected Qwen2.5-7B as the implementation solution for the patient agent. For a more detailed description of the dialogue flow, please refer to Supplementary Material.